# SurgicAI: A Fine-grained Platform for Data Collection and Benchmarking in Surgical Policy Learning

### v1

@@ 4.2 Teleoperation and Dataset Collection
Our environment inherits the benefit of SRC to support various user interfaces for controlling the PSM arms, including graphical user interfaces and devices such as Geomagic and Razer Hydra [37]. Meanwhile, SurgicAI includes unified data processing that transform this raw data into a format compatible with the Gymnasium API. This conversion process generates well-structured transitions involving basic observations, actions, rewards, and terminal states, thereby ensuring the data’s consistency and usability for training policies. This flexibility allows users to switch between different devices for data collection without requiring any modifications, thereby enhancing the efficiency of the data collection process. Additionally, beyond providing a platform for data collection, SurgicAI also offers access to some of the collected trajectories for training purposes, including around 30 human trajectories and 100 heuristic trajectories. These are available on our project website for further use.
@@ 5 Implementation and Experiment Results
This section covers the implementation of algorithms and experiments, including data analysis. Empirically, we aim to demonstrate the following:
@@
    The performance of different RL and IL algorithms in suturing tasks, especially the importance of imitation learning, particularly behavior cloning, for successful policy development.
@@
Performance of RL and IL algorithms: To evaluate the performance of various algorithms, we tested Proximal Policy Optimization (PPO) [40], Deep Deterministic Policy Gradient (DDPG) [41], Soft Actor-Critic (SAC) [42], and Twin Delayed DDPG (TD3) [43], along with ablation tests involving Hindsight Experience Replay (HER) [39] and Behavior Cloning (BC) [44]. These tests employed sparse rewards, circumventing the need for reward shaping. Detailed environment settings and network hyperparameters are provided in Sections 3.2 and A.1, respectively. By default, the BC loss weight during testing is set to 0.5 and we further discuss the impact of the weight in Section A.4.
Table 1: Mean success rates (expressed as probabilities) for selected subtasks using different learning algorithms after 150k training steps. Each result is evaluated over 100 episodes.
@@
Algorithm Approaching Placing Inserting Handoff Pullout
PPO 0.00 0.00 0.00 0.00 0.00
DDPG 0.00 0.01 0.00 0.00 0.00
SAC 0.01 0.02 0.00 0.00 0.00
TD3 0.02 0.01 0.01 0.01 0.01
BC 0.95 1.00 0.98 0.96 0.97
TD3+HER 0.10 0.12 0.08 0.09 0.10
TD3+BC 0.80 0.85 0.78 0.80 0.82
TD3+HER+BC 0.90 0.90 0.88 0.85 0.88
Table 1 summarizes the success rates for each RL algorithm. Pure RL methods exhibited nearly zero success rates across all subtasks, highlighting the challenges in exploration under sparse reward conditions. Even with techniques like HER, performance improvements remained limited due to the curse of dimensionality and the demanding precision requirements. Conversely, integrating BC with RL significantly enhanced performance, with TD3+HER+BC yielding substantial success rates. Notably, BC outperformed other methods by providing a robust initial policy and mitigating exploration challenges. Moreover, RL-based methods encountered difficulties with specific tasks such as Inserting and Handoff, which are prone to errors. For instance, in the Inserting task, agents often opted for strategies with minimal-step completion to reduce reward discounting, occasionally leading to suboptimal trajectories and missed needle exit points. The Handoff task demands precise gripper manipulation to grasp a small needle segment; any collision with the phantom can alter the gripper’s pose and result in failure. In these contexts, BC demonstrated superior performance by adhering more closely to expert trajectories.
Table 2: Mean MSE loss for each subtask using image-based imitation learning after 30 epochs. Data collection and training are performed from two camera views (front and back).
Task Name Camera View R3M CLIP ImNet
Approaching Front 0.394 0.401 0.458
Back 0.398 0.407 0.469
Placing Front 0.379 0.355 0.402
Back 0.377 0.367 0.417
Inserting Front 0.362 0.368 0.397
Back 0.370 0.365 0.400
Handoff Front 0.396 0.407 0.460
Back 0.398 0.412 0.471
Pullout Front 0.313 0.320 0.356
Back 0.319 0.324 0.362
Effectiveness of Visual Representaions In our study of image-based imitation learning, we evaluated the mean MSE loss across various subtasks using different visual representations, as shown in Table 2. Detailed settings can be found in Section A.2. We chose MSE loss as the metric over success rate because it provides a quantitative measure of the models’ continuous performance, crucial for assessing consistency with the demonstrated trajectory. The results evaluate the performance of three state-of-the-art visual representations widely used in robotics: Contrastive Language-Image Pretraining (CLIP) [45], Pretraining Reusable Representations for Robot Manipulation (R3M) [46], and Pretrained ResNet-50 [47] using the ImageNet dataset (ImNet) [48] across two camera views (front and back). The lowest MSE loss, consistently observed in the Pullout task from the front camera view, suggests that certain tasks and perspectives provide more stable conditions for these learning algorithms. Notably, R3M slightly outperforms CLIP and ImNet in tasks requiring precise spatial alignment, such as Inserting and Pullout. This nuanced visualization aids in understanding the comparative performance at a glance and underscores the impact of camera perspective on the efficacy of image-based imitation learning. These insights are crucial for refining task-specific training strategies and future enhancements of robotic learning frameworks.
@@
Hierarchical Multi-stage Suturing Performances: While testing the final points, in the configuration of our single-policy approach, we utilized a sparse reward structure, awarding the agent upon the successful completion of each subtask. The detailed result is shown in Table 3. When BC was integrated, expert demonstrations encompassing the entire suturing sequence were employed to guide the agent’s actions. Although BC effectively managed individual subtasks, its performance was limited in longer task sequences. This limitation primarily stemmed from distribution shifts, leading to a significant decline in success rates as task complexity escalated. In contrast, our hierarchical task decomposition framework demonstrated substantial performance enhancements, achieving over a $50\%$ success rate in the complete suturing procedure. This marked improvement highlights the framework’s robustness and efficacy in managing complex surgical tasks.
@@ 6 Conclusions, Limitations, and Future Plans
In conclusion, SurgicAI represents a significant milestone by providing a comprehensive and standardized pipeline for suturing automation, implementing a robust hierarchical learning framework, and creating a diverse suite of manipulative tasks that benchmark various learning approaches. The results demonstrate that our framework significantly enhances the suturing performance, underscoring its practical impact on surgical automation. The remaining challenges and future work are as follows:
More realistic simulations: SurgicAI currently utilizes the simulator version from the 2022-2023 AccelNet Surgical Robotics Challenge [49]. As the challenge evolves, incorporating new models of real phantoms, we will update SurgicAI accordingly. Our datasets and pipelines will be adapted to ensure compatibility with different versions of phantoms and instruments. Our future work also includes developing a more realistic simulator with higher sampling rate, higher rendering quality and more precise modeling of medical instruments to reduce the domain gap in Sim2Real transfer, ultimately facilitating the fine-tuning and transfer of our policies to real-world suturing.
More algorithms and tasks: SurgicAI has been tested with only a limited number of algorithms. We encourage the open-source community to contribute to expanding the range of algorithms tested with our system, enhancing its robustness and versatility. In the future, we plan to explore the integration of vision-language models to improve high-level policy decision-making. We also plan to expand our subtask library and enhance the system’s robustness against more complex and unexpected scenarios. With more medical simulation environments based on AMBF [50, 51, 52], we see significant potential to deploy SurgicAI in various surgical scenarios.
Collaboration The surgical robotics challenge, published as an annual competition [49], invites teams from around the world to contribute their trajectories and approaches to suturing automation. To foster a collaborative community where participants can advance algorithms and augment datasets, we aim to further advance SurgicAI via more realistic simulations and more diverse algorithms and tasks and look forward to collaborating with the broader research community to realize these goals.
@@ Acknowledgments
This work was supported by NSF AccelNet award OISE-192735 and an agreement between Johns Hopkins University and Medical Robotics Center (MRC). AL is partially supported by the Discovery Award of the Johns Hopkins University and a seed grant from JHU Institute of Assured Autonomy.
@@ A.1 Reinforcement Learning Settings
We perform all our training process on one Tesla T4 GPU hosted on Microsoft Azure. All sampling during the training are running on headless mode. We implement and modify RL algorithms from SB3 Library. Both the HLP and LLPs utilize an actor-critic architecture, with the actor and critic networks comprising four fully-connected layers, each containing 256 hidden units and ReLU activations. The LLPs’ actions are scaled to the range [-1, 1] using a Tanh activation function in the actor network. Networks are trained using the ADAM optimizer and a default discount factor of 0.99. All LLPs are trained independently in separate Gymnasium environments. The HLP is trained afterward, using the fixed policies of the well-trained LLPs. Besides, When HER is applied for LLP, three additional goals are sampled per episode to enhance the training data, using the ’future’ strategy for goal selection. The demonstration data is collecting from heuristic trajectories. In our experimental setup, the HLP employs PPO algorithm. Meanwhile, we test various RL algorithms for the LLPs.

### v3

@@ 4.2 Teleoperation and Dataset Collection
Our environment inherits the benefit of SRC to support various user interfaces for controlling the PSM arms, including graphical user interfaces and devices such as Geomagic and Razer Hydra [18]. Meanwhile, SurgicAI incorporates a standardized data processing pipeline that converts raw data into a format compatible with the Gymnasium API. This process organizes observations, actions, rewards, and terminal states into a structured transition format. This flexibility allows users to switch between different devices for data collection without requiring any modifications, thereby enhancing the efficiency of the data collection process. Additionally, beyond providing a platform for data collection, SurgicAI also offers access to some of the collected trajectories for training purposes, including around 30 human trajectories and 50 heuristic trajectories. These are available on our project website [^2] https://github.com/surgical-robotics-ai/SurgicAI/tree/main/RL/Expert_traj for further use.
@@ 5 Implementation and Experiment Results
This section covers the implementation of algorithms and experiments, including result analysis. Empirically, we aim to demonstrate the following:
@@
    The performance of different RL and IL algorithms in suturing tasks, especially the importance of expert demonstration data for successful policy development.
@@
Evaluation Metrics: In this study, we utilized three primary evaluation metrics to assess the performance of low-level policies:
  1. 1.
    Success Rate: The proportion of episodes where the task was successfully completed out of the total number of episodes. This metric reflects the algorithm’s ability to perform the task successfully.
  2. 2.
    Trajectory Length: The cumulative distance, measured in millimeters, covered by delta actions during a successful episode. This metric provides insight into the physical efficiency of task completion.
  3. 3.
    Time Steps: The total number of discrete actions taken to complete the task. This metric measures temporal efficiency, indicating how quickly the agent completes the task.
These metrics collectively help evaluate the effectiveness of the low-level policies, where the ideal outcome is a high success rate, minimal trajectory length, and fewer time steps.
@@
Performance of RL and IL algorithms: To evaluate the performance of various algorithms, we tested online algorithms like Proximal Policy Optimization (PPO) [43], Deep Deterministic Policy Gradient (DDPG) [44], Soft Actor-Critic (SAC) [45], and Twin Delayed DDPG (TD3) [46], along with ablation tests involving Hindsight Experience Replay (HER) [42] and Behavior Cloning (BC) [47], as well as offline algorithms including Advantage Weighted Actor-Critic (AWAC) [48], Batch-Constrained Q-learning (BCQ) [49], Implicit Q-learning (IQL) [50], and Calibration Q-learning (CalQL) [51]. Both sparse and dense rewards are applied for evaluation. Detailed environment settings and network hyperparameters are provided in Sections 3.2 and A.1, respectively. By default, the BC loss weight during testing is set to 0.5 and we further discuss the impact of the weight in Section A.5.
Table 1: Performance results of different RL algorithms after 150k training steps, with results averaged over 20 episodes using 5 random seeds. ’/’ indicates that no successful episodes were collected.
@@
Algorithm Reward Type Criteria Grasping Placing Inserting Handoff Pullout
PPO Dense Success Rate 0.22 ± 0.10 / / 0.20 ± 0.20 0.40 ± 0.49
Trajectory Length (mm) 31.80 ± 3.84 / / 58.54 ± 2.22 43.59 ± 1.03
Time step 81.00 ± 8.05 / / 97.09 ± 4.51 85.21 ± 4.23
Sparse Success Rate / / / / /
DDPG Dense Success Rate 0.02 ± 0.04 / / / 0.04 ± 0.04
Trajectory Length (mm) 38.72 ± 2.56 / / / 40.53 ± 1.56
Time step 87.50 ± 9.50 / / / 83.94 ± 3.64
Sparse Success Rate / / / / /
SAC Dense Success Rate / / / / /
Sparse Success Rate / / / / /
BC / Success Rate 0.96 ± 0.05 1.00 ± 0.00 0.97 ± 0.02 0.83 ± 0.24 0.92 ± 0.09
Trajectory Length (mm) 51.31 ± 14.60 66.35 ± 20.36 42.43 ± 2.51 64.65 ± 3.21 49.88 ± 1.74
Time step 100.15 ± 26.02 120.87 ± 26.72 89.80 ± 3.98 113.60 ± 5.15 95.80 ± 3.60
TD3 Sparse Success Rate / / / / /
TD3+HER Sparse Success Rate 0.11 ± 0.12 0.15 ± 0.09 0.05 ± 0.05 0.12 ± 0.09 0.15 ± 0.14
Trajectory Length (mm) 46.55 ± 7.49 55.26 ± 11.25 38.71 ± 1.88 58.32 ± 10.23 45.23 ± 2.46
Time step 86.18 ± 21.85 108.62 ± 23.88 79.28 ± 3.21 95.43 ± 8.32 90.46 ± 5.65
TD3+BC Sparse Success Rate 0.85 ± 0.10 0.88 ± 0.08 0.82 ± 0.15 0.79 ± 0.14 0.90 ± 0.10
Trajectory Length (mm) 52.45 ± 15.90 63.39 ± 15.60 46.73 ± 4.59 65.98 ± 3.92 48.52 ± 1.85
Time step 100.07 ± 18.10 116.71 ± 32.88 87.00 ± 8.89 108.80 ± 4.83 94.00 ± 3.90
TD3+HER+BC Sparse Success Rate 0.96 ± 0.06 0.97 ± 0.09 0.91 ± 0.07 0.98 ± 0.02 1.00 ± 0.00
Trajectory Length (mm) 47.89 ± 19.25 59.27 ± 19.92 41.74 ± 2.82 61.61 ± 6.15 41.88 ± 2.49
Time step 89.61 ± 25.36 110.63 ± 31.31 83.25 ± 3.30 105.00 ± 4.30 88.20 ± 4.93
AWAC Dense Success Rate / / 0.96 ± 0.06 0.98 ± 0.09 /
Trajectory Length (mm) / / 44.06 ± 4.23 63.40 ± 5.96 /
Time step / / 89.32 ± 2.32 94.00 ± 10.32 /
BCQ Dense Success Rate 0.94 ± 0.10 0.92 ± 0.09 0.95 ± 0.08 0.96 ± 0.04 0.93 ± 0.10
Trajectory Length (mm) 49.22 ± 14.36 61.73 ± 9.10 42.38 ± 2.31 69.78 ± 4.63 45.02 ± 3.69
Time step 94.75 ± 19.36 117.90 ± 24.46 93.00 ± 4.52 106.00 ± 7.50 93.23 ± 3.21
IQL Dense Success Rate 0.95 ± 0.02 0.91 ± 0.09 0.94 ± 0.05 0.95 ± 0.06 0.95 ± 0.08
Trajectory Length (mm) 47.40 ± 15.13 67.28 ± 11.52 43.46 ± 2.51 72.21 ± 3.43 44.28 ± 2.34
Time step 92.15 ± 21.57 121.85 ± 23.56 48.00 ± 4.12 110.34 ± 6.21 95.21 ± 4.31
CalQL Dense Success Rate / / / / /
Table 1 presents the results of each RL algorithm. Pure online RL methods struggled to achieve significant success rates, especially under sparse reward settings. For example, both PPO and DDPG exhibited near-zero success rates for most subtasks, underscoring the challenge of exploration in sparse reward environments. Even with techniques like HER, improvements were limited due to the curse of dimensionality and stringent precision requirements.
Performance improved slightly with dense rewards but often led to suboptimal trajectories, as agents tended to favor the shortest path to task completion. Specific tasks like inserting and handoff posed significant challenges. In the inserting task, agents often adopt strategies with minimal-step completion to minimize reward discounting, occasionally resulting in suboptimal trajectories and missed needle exit points. The handoff task requires precise gripper manipulation to grasp a small needle segment; any collision with the phantom can alter the gripper’s pose and lead to failure.
In contrast, integrating BC with RL demonstrated superior performance by closely adhering to expert trajectories. TD3+HER+BC, as demonstrated in A.4, achieved high success rates. While BC alone could attain comparable success rates, the addition of the RL loss optimized the policy, resulting in more efficient trajectories and shorter completion times. For example, in the approaching task, TD3+HER+BC reduced the trajectory length by approximately 5 mm and decreased the completion time by 10 steps compared to BC alone, demonstrating improved efficiency in both spatial and temporal dimensions.
Offline algorithms generally outperformed online ones, particularly in dense reward settings. For instance, BCQ consistently achieved high success rates across all tasks, including 0.95 for the inserting task and 0.96 for handoff. However, algorithms, such as AWAC and CalQL, struggled on tasks like approaching and placing, achieving zero success rates during evaluation. This is because their reliance on accurate advantage or Q-value estimates made them vulnerable in tasks where the offline data was not comprehensive enough to support reliable learning. This performance could be improved with more diverse and abundant offline data.
Table 2: Performance results for each subtask using image-based imitation learning after 30 epochs. Each visual representation was tested under two camera views, with results averaged over 20 episodes using 5 random seeds.
Task Criteria R3M CLIP ImNet RandomNet
front back front back front back front back
Success Rate 0.22±0.14 0.35±0.04 0.22±0.06 0.32±0.06 0.08±0.06 0.10±0.00 0.03±0.05 0.03±0.02
Trajectory length (mm) 40.81±6.44 44.04±7.63 50.91±6.03 46.32±8.65 47.88±8.22 49.59±10.04 54.85±15.33 41.60±4.11
Grasp Time cost (step) 93.18±19.59 94.92±15.75 109.63±14.20 97.65±22.61 89.00±20.66 91.83±18.38 102.00±31.00 74.50±8.50
Success Rate 0.40±0.07 0.22±0.06 0.30±0.07 0.28±0.09 0.27±0.12 0.20±0.18 0.08±0.06 0.02±0.02
Trajectory length (mm) 60.80±9.94 60.06±8.24 70.23±11.43 71.89±7.77 67.80±9.46 77.31±12.60 80.70±0.00 69.02±2.90
Place Time cost (step) 125.88±17.97 134.55±32.86 171.60±12.43 165.94±29.86 181.46±13.32 160.46±20.12 131.00±0.00 120.20±4.83
Success Rate 0.80±0.04 0.98±0.02 0.62±0.06 0.78±0.02 0.77±0.06 0.83±0.02 0.03±0.02 0.07±0.03
Trajectory length (mm) 56.12±3.72 54.09±2.88 59.34±4.08 58.42±3.90 66.70±6.31 64.52±3.60 71.90±3.61 69.02±2.90
Handoff Time cost (step) 107.98±7.09 116.92±5.41 125.65±15.16 118.47±6.23 122.48±17.32 117.68±8.78 147.50±8.50 110.20±4.83
Success Rate 0.82±0.12 0.63±0.09 0.62±0.02 0.60±0.04 0.35±0.04 0.20±0.07 0.10±0.00 0.08±0.05
Trajectory length (mm) 44.86±2.64 50.20±3.40 49.42±2.35 52.43±2.38 50.59±2.87 54.10±4.77 49.98±1.25 46.96±1.39
Pullout Time cost (step) 76.00 ± 5.58 83.47±4.31 83.92±3.11 94.72±5.17 74.95±6.78 81.71±12.43 78.00±2.58 74.25±1.92
Effectiveness of Visual Representations: In our study of image-based imitation learning, we applied the same evaluation metrics to assess different visual representations, as shown in Table 2. Detailed settings are provided in Section A.2. We evaluated the performance of three popular visual representations: Contrastive Language-Image Pretraining (CLIP) [52], Pretraining Reusable Representations for Robot Manipulation (R3M) [53], and a pretrained ResNet-50 model [54] using the ImageNet dataset (ImNet) [55]. Also, we included a randomly initialized ResNet-50 (RandomNet) as a baseline for comparison. Each model was tested with two camera views: front and back.
A high success rate was consistently observed in the handoff task using the back camera view, suggesting that certain tasks and perspectives provide more favorable conditions for these learning algorithms. R3M slightly outperformed CLIP and ImNet in tasks requiring precise spatial alignment, such as handoff and pullout. Pretrained models consistently outperformed RandomNet in terms of success rate. This is likely because they were trained on large, diverse datasets, enabling them to learn valuable representations such as semantic information and image patterns. These pretrained features can be fine-tuned or directly applied to new domains, significantly reducing computational costs and sample complexity compared to training an encoder from scratch, which can lead to overfitting given the limited scale of our data.
@@
Hierarchical Multi-stage Suturing Performances: While testing the final points, in the configuration of our single-policy approach, we utilized a sparse reward structure, awarding the agent upon the successful completion of each subtask. The detailed result is shown in Table 3. When BC was integrated, expert demonstrations encompassing the entire suturing sequence were employed to guide the agent’s actions. Although BC effectively managed individual subtasks, its performance was limited in longer task sequences. This limitation primarily stems from the accumulation of errors over extended execution, which progressively leads the agent into states that differ substantially from those represented in the demonstration data. Consequently, a distribution shift occurs between the training data and the actual states encountered during the suturing process. In contrast, our hierarchical task decomposition framework demonstrated substantial performance enhancements, achieving over a $50\%$ success rate in the complete suturing procedure. This marked improvement highlights the framework’s robustness and efficacy in managing complex surgical tasks.
@@ 6 Conclusions, Limitations, and Future Plans
In conclusion, SurgicAI provides a comprehensive and standardized pipeline for suturing automation, implementing a robust hierarchical learning framework, and creating a diverse suite of manipulative tasks that benchmark various learning approaches. The results demonstrate that our framework significantly enhances the suturing performance, underscoring its practical impact on surgical automation. The remaining challenges and future work are as follows:
More realistic simulations: SurgicAI currently utilizes the simulator version from the 2022-2023 AccelNet Surgical Robotics Challenge [56]. As the challenge evolves, incorporating new models of real phantoms, we will update SurgicAI accordingly. Our datasets and pipelines will be adapted to ensure compatibility with different versions of phantoms and instruments. We are also developing a more realistic rendering pipeline [57] to reduce the domain gap in Sim2Real transfer, as well as injecting the kinematic error [58] of physical robots, ultimately facilitating the fine-tuning and transfer of our policies to real-world suturing.
More algorithms and tasks: SurgicAI has been tested with a limited set of algorithms, but still the image-based policies exhibit suboptimal performance, particularly in terms of success rate and stability. The performance also fluctuates significantly with changes in camera perspectives, suggesting that dynamic-view and wrist-mounted cameras could help capture more comprehensive information. To address these issues, we encourage further efforts to improve the robustness of image-based policies. Besides, our current policies rely on Markovian single-step methods, which are inadequate for tasks involving temporal dependencies like pauses or corrections [59]. To overcome this, incorporating advanced algorithms such as the Action Chunking Transformer (ACT) [59] and Diffusion Policy (DP) [60] will be essential. Looking ahead, we aim to integrate vision-language models to improve high-level decision-making, expand our subtask library, and strengthen the system’s performance in complex and unexpected scenarios. With more medical simulation environments and applications based on AMBF [61, 62, 63, 64, 65], SurgicAI has the potential for deployment across various surgical applications.
Collaboration: The surgical robotics challenge, published as an annual competition [56], invites teams from around the world to contribute their trajectories and approaches to suturing automation. To foster a collaborative community where participants can advance algorithms and augment datasets, we aim to further advance SurgicAI via more realistic simulations and more diverse algorithms and tasks. We look forward to collaborating with the broader research community to realize these goals.
@@ Acknowledgments
This work was supported by NSF AccelNet award OISE-1927354 and an agreement between Johns Hopkins University and Medical Robotics Center (MRC). AM is partially supported by Discovery Award titled Simulation Assisted Navigation for Skull-base Surgery by Johns Hopkins University (JHU). AL is partially supported by the Amazon Research Award, the Discovery Award of the JHU and a seed grant from the JHU Institute of Assured Autonomy. The full team for the development of the underlying SRC simulation and other infrastructures, on which we build our framework, includes Juan Antonio Barragan Noguera, Hisashi Ishida, Haoying Zhou and Dr. Adnan Munawar.
@@ A.1 Reinforcement Learning Settings
We perform all our training processes on one Tesla T4 GPU hosted on Microsoft Azure. All sampling during the training are running in headless mode. We implement and modify RL algorithms from the SB3 Library. Both the HLP and LLPs utilize an actor-critic architecture, with the actor and critic networks comprising four fully-connected layers, each containing 256 hidden units and ReLU activations. The LLPs’ actions are scaled to the range [-1, 1] using a Tanh activation function in the actor network. Networks are trained using the ADAM optimizer and a default discount factor of 0.99. All LLPs are trained independently in separate Gymnasium environments. The HLP is trained afterward, using the fixed policies of the well-trained LLPs. Besides, When HER is applied for LLP, three additional goals are sampled per episode to enhance the training data, using the ’future’ strategy for goal selection. The demonstration data is collected from heuristic trajectories. In our experimental setup, the HLP employs a PPO algorithm. Meanwhile, we test various RL algorithms for the LLPs.